OpenStudy (jazzyfa30):

convert each angle in degrees to radians. A)125 degrees B)-51 degrees C)17 degrees

OpenStudy (jazzyfa30):

@johnweldon1993

OpenStudy (ipwnbunnies):

To convert degrees to radians, multiply by pi/180. For example: \[180 degrres*\frac{\pi}{180} = \pi radians\]